# Self interest cumulative subtraction games

### v1

@@ 1 Introduction
We study two-player alternating play perfect information games, and we adapt the following definition of a self-interest symmetric cumulative subtraction game $S\subset\mathbb{N}=\{1,2,\ldots\}$. There is a single heap of pebbles, of size $h\in\mathbb{N}_{0}$, and the players alternately remove $s\in S$ pebbles from the heap, until the current player is not able to move from some position $h$ because, for all $s\in S,h-s<0$. The players are called Alice and Bob, and whatever they remove, from the common heap, they put in their own individual pocket. Each player aims to maximize the size of their pocket (i.e., their cumulative utility) when the game ends.
@@
We adapt standard tools from game theory, in particular the classical backward induction procedure, applied to player profiles in Pure Subgame Perfect Equilibrium (PSPE). To avoid ambiguity in the recursive evaluation and to ensure the game tree is well-defined, we assume deterministic tie-breaking in cases of indifference. We refer to a pair of tie-breaking conventions, one for each player, as deterministic if they satisfy six properties. They cannot be random or hidden in any way, and they must be universal, meaning that they are applied to every position whenever there is indifference between options, with respect to PSPE-computation. They must be uniform, meaning that the same conventions are used across all positions. They must be Markovian (or history-independent), that is, they cannot depend on previous actions in any way. Finally, they are unilateral, meaning that each player’s convention is independent of the other player’s choice. One can prove by induction that PSPE play is indeed well defined by using these tie-breaking axioms.
@@
By using a pair of deteministic tie-breaking conventions, players still maximize their own utility. However, when multiple options yield the same personal outcome, the choice among them may influence the opponent’s utility. We classify this influence as friendly when it benefits the opponent and antagonistic when it harms them. This effect is understood to be local, meaning that it depends only on outcome discrepancies within any given pair of deterministic tie-breaking conventions. As shown in the Appendix, page Appendix: Unilateral deviation for larger subtraction sets, such local choices can have globally counterintuitive consequences, when viewed across ranges of possible conventions.
@@
Deterministic tie-breaking conventions are commitments that are applied consistently throughout the game tree, rather than as decisions made at individual nodes. For instance, given a pair of tie-breaking conventions, for every heap size, even one as large as $10^{10}$, a player is required to know whether they are indifferent. If so, for example a player who adopts a friendly convention, must then select a move that maximizes the opponent’s result, based on the already computed PSPE utilities of the resulting subpositions.
@@
Following standard assumptions in combinatorial game theory, we assume the existence of such perfectly rational and computationally capable players. When both players adopt the friendly convention, we refer to the resulting setting as FvF or the friendly convention. If both adopt antagonistic conventions, we denote it AvA or the antagonistic convention. These will serve as our primary self-interest frameworks throughout the paper. [^1] In our model, the tie-breaking rule is fixed in advance and is not determined as part of the PSPE. Analyzing strategic choice over tie-breaking behavior would require a different model, such as a simultaneous-play normal form with mixed strategies. Here, we focus instead on the recursive structure and properties of combinatorial games.
@@
In an antagonistic self-interest cumulative subtraction game, we define the (self-interest) outcome function of a game $S\subset\mathbb{N}$ for the two players as the ordered pair $o_{\scriptscriptstyle{\rm AvA}}(h)=(o^{1}_{\scriptscriptstyle{\rm AvA}}(h),o^{2}_{\scriptscriptstyle{\rm AvA}}(h))$, where $h\in\mathbb{N}_{0}=\mathbb{N}\setminus\{0\}$ represents the current heap size. The cumulative utilities (a.k.a individual outcomes) can be recursively calculated using the outcome of heaps of smaller size. Let the first player choose $s\in S$ as their first move in PSPE play. Then $o^{1}(h)=s+o^{2}(h-s)$ and $o^{2}(h)=o^{1}(h-s)$. Thus, after the first move the heap size is reduced to (hs)(h-s) and the order of the players is switched (since the next move is made by the originally second player). Here, whenever there is indifference (i.e., more than one action leading to the same PSPE cumulation for the current player), the action chosen minimizes the PSPE cumulation for the other player, representing antagonistic behavior. In the case of friendly cumulative play, the individual outcomes $o^{1}_{\scriptscriptstyle{\rm FvF}}$ and $o^{2}_{\scriptscriptstyle{\rm FvF}}$ are analogous, but the action chosen maximizes the other player’s cumulation in case of indifference, representing friendly behavior. Note that these two settings remain symmetric, in the sense that the outcome pairs do not depend on “who starts”, Alice or Bob. Thus, the setting represents a distant relative to the classical ‘impartial’ combinatorial games.
The first question is whether there will be any outcome-discrepancies between both players adapting friendly tie-breaking rules versus both players adapting antagonistic tie-breaking rules. Let player $i$’s discrepancy be $\delta^{i}(h)=o^{i}_{\scriptscriptstyle{\rm AvA}}(h)-o^{i}_{\scriptscriptstyle{\rm FvF}}(h)$, for $i\in\{1,2\}$. In Table 1 we exemplify this result with the ruleset $S=\{3,5\}$ for some initial heap sizes.
@@
$\delta^{2}(x)$ 0 0 0 0 0 0 0 0 0 0 0 0 0 0 1-1 0
The corresponding PSPE move sequence is 3-3-5-3, where Alice makes the first and third removal. We experimentally discovered a number of such examples of discrepancies in the two and three action settings, summarized in Section 4, with related conjectures. The main result of this paper, Theorem 2, is that, if the subtraction set has size 2, then, for any heap size, each player’s utility is no worse in the FvF setting than the AvA setting. We conjecture that this result continues to hold if $|S|>2$; see Section 2.
Readers familiar with zero-sum games may note that the move sequence in the previous paragraph cannot appear in an optimal-play zero-sum setting, where player removals add and subtract to a common score, respectively. Alice would instead start by removing 5, and hence Bob must take 5, and Alice gets the last move by taking 3. This, of course leads to a loss in “social welfare” as the whole heap would not be consumed. It related topic explores the similarities between the $\rm AvA$} antagonistic and the zero-sum settings. We explore discrepancies between the games in a follow up paper [28], where we prove that if $|S|=2$, stated here as a conjecture (see Section 5), then optimal play coincides for zero-sum games and the antagonistic setting. However, experimental studies reveal that discrepancies appear already for $|S|=3$ (related conjectures appear in Section 5).
Let us return to the main topic for this paper, concerning discrepancies between friendly and antagonistic tie-breaking rules. If the players play many rounds of the same game, and they have so far both been committed to the friendly setting, could either player benefit by deviating and becoming antagonistic?
@@
Could any personalized player benefit by a unilateral deviation from a deterministic friendly tie-breaking?
The Penguin narrative will be revisited in the Appendix. Let us outline the content of this paper. In Section 2 we generalize the outcome function to four deterministic tie-breaking pairs and explain the road map to the main theorem for two-action rulesets. In Section 3, we prove the main theorem, Theorem 2, via a few lemmas, highlighting first player advantage, greedy play vs. sacrifice, tie-breaking monotonicity and more. In Section 4, we outline a follow up project with several structurally detailed conjectures concerning FvF vs AvA (those are the conjectures that motivated this study), while in Section 5, we outline some conjectures by instead comparing AvA with zero-sum play. In Section 6 we discuss further open problems, and in Section 7, we mention some related work. Finally, in the Appendix (page Appendix: Unilateral deviation for larger subtraction sets), we illustrate how cases where $|S|>2$ can differ from two action rulesets, in globally counter intuitive ways.
2 The model and main result
Question 1 induces two more settings. Let us list them all.
@@ 2 The model and main result
    Alice, as a starting player, acts friendly and Bob is antagonistic, in case of indifference;
@@
    Bob, as the second player, acts friendly whereas Alice is antagonistic, in case of indifference;
@@
Observe that this requires a generalization of the outcome function. Let $\tau=\{{\rm FvF},{\rm AvF},{\rm FvA},{\rm AvA}\}$, and let $X\in\tau$. The dual of $X$, denoted $\mathsf{d}(X)$, is defined by:
@@
Given a tiebreaking convention, we usually refer to a pair of (EGT) utilities in PSPE simply as a (CGT) outcome. Note that even if the ruleset is symmetric, since the players may have different tie-breaking rules, a game played is not necessarily symmetric in the strict sense. If one player is antagonistic at some stage of play then they are committed to being antagonistic at all stages of play, and similarly for “friendly”. It is a situation that cannot appear in standard CGT impartial games. The move options, in this sense, depend on their committed tie-breaking rule. To our best knowledge such situations have not yet been studied in the literature.
Thus there are four deviation cases to consider, listed as FvF $\rightarrow$ AvF, FvF $\rightarrow$ FvA, AvA $\rightarrow$ FvA and AvA $\rightarrow$ AvF. For example in FvF $\rightarrow$ AvF, in the FvF setting both players resolve ties in a friendly manner, while in AvF, Alice (playing first) deviates with an antagonistic tie-breaking rule.
In the case $|S|=2$, we here prove that a unilateral deviation from friendly tie-breaking to antagonistic ditto never benefits either player; as we will see, whenever starting from FvF, the deviating player will be unaffected, while the other player may suffer in some cases. As a consequence we will get:
@@
Consider any subtraction set $S$, with $|S|=2$, and suppose both players act friendly in case of indifference. Then each player’s PSPE utility is never worse than if both players have antagonistic tie-breaking rules.
Our methods to prove this result do not apply in general if $|S|>2$, and we provide a counter example, when unilateral deviation is non-monotone, namely if $S=\{4,5,9\}$, in Table 3 in the Appendix. We will study the two unilateral deviations from FvF, to FvA and AvF, while the converse/symmetric deviations from AvA, to AvF and FvA are analogous. A unilateral deviation from friendly to antagonistic may, in rare cases, be beneficial for the deviating player; still, we have numerical evidence that Theorem 2 continues to hold for any subtraction set of size greater than 2 (see Section 6). However, we currently have no formal method to jump directly from both friendly to both antagonistic.
@@
Consider ordered pairs of integers, $(a,b)$ and $(c,d)$. Then $(a,b)\leq(c,d)$ if $a\leq c$ and $b\leq d$, and similar with $\geq$”.
3 Unilateral tie-breaking deviation and the proof of main result
Suppose both players follow deterministic tie-breaking conventions. We are interested in comparing the outcome under one such setting, where both players are universally friendly, with the setting in which exactly one of the players deviates, by committing to an antagonistic tie-breaking rule.
Whenever the subtraction set, $S=\{s_{2},s_{1}\}$, has exactly two actions, then we claim that the deviating player’s PSPE-utilities will be exactly the same as if they would have remained friendly, and the other player’s PSPE-utilities cannot improve. This claim holds independently of who starts, and we prove it in Lemma 7 in this section.
In the coming sections, we will fix the convention $s_{2}<s_{1}$.
Through Lemma 7, we see how the claims in the first paragraph combine in an induction proof. For any two-action subtraction set $S$, let $\Delta=s_{1}-s_{2}>0$. A greedy move from $h\geq s_{1}$ is to play to $h-s_{1}$, and otherwise a player sacrifices. The endgame refers to play to a heap size $h<s_{1}$. The slack is a leftover that cannot be attained by either player, and so it refers to a heap of size $h<s_{2}$.
There is a general proof of the next result in the zero-sum setting [14]. Here it only holds in the two-action setting. See Table 3 in the Appendix on page Appendix: Unilateral deviation for larger subtraction sets, where $S=\{4,5,9\}$ and where heaps of sizes $61,75,\ldots$ have strictly better PSPE utility for the second player.
@@ 3 Unilateral tie-breaking deviation and the proof of main result
Recall $s_{1}>s_{2}$. Suppose Alice plays greedily, until she cannot play $s_{1}$. Then she gets utility $ks_{1}+\ell s_{2}$, some $k>0$, some $\ell\geq 0$. If Bob gets to play $k$ copies of $s_{1}$ then Alice has first player advantage at the endgame too, and hence, Bob cannot get more than $ks_{1}+\ell s_{2}$. If Bob can only collect k1k-1 copies of $s_{1}$, then he can at most get one more copy of $s_{2}$ than Alice, but their utility difference is at least $\Delta>0$ in Alice’s favor. ∎
@@
The usefulness of a sacrifice may appear limited while playing on a single heap of tokens. However, simple arithmetic shows that if $S=\{2,3\}$ then Alice benefits strictly when playing from a heap of size $7$, and if $S=\{3,4\}$, then she can benefit strictly by sacrificing twice, playing from $h=17$. There are many more such examples, and those are analogous to the zero-sum setting studied in [14], but in two-action zero-sum games, the usefulness of a sacrifice is limited to basic aritmetic considerations for small heap sizes. However, self-interest play appears more multifaceted. As shown in Table 1, there are situations where, under a deterministic tie-breaking rule, Alice sacrifices solely to benefit Bob. Note that the example occurs in the non-dominant regime. By contrast, the dominant regime generally exhibits less complex behavior.
@@
The proof will use a strategy-stealing argument to establish a contradiction. Suppose Alice sacrifices in her first move and profits strictly in PSPE play. This means that she must get at least $k$ more $s_{2}$ actions to compensate for her initial sacrifice, if $s_{1}=ks_{2}+r$, $0\leq r<k$. However, since $k\geq 2$, this requires that Bob also plays a sacrifice at some point. For otherwise Bob gets at least $3\Delta-s_{1}\geq\Delta>0$ larger utility than Alice, while Lemma 4 assures that Alice as a starting player never gets worse than Bob in PSPE play. But Bob can replace this sacrifice with $s_{1}$, by stealing $ks_{2}+r\leq(k+1)s_{2}$ of Alice’s required removal segment. Hence Alice’s first sacrifice cannot lead to a strict gain. ∎
A consequence of this result is that, in case of dominance, $s_{1}$ is never a bad choice for the current player (if feasible). And indeed, experimental results do point towards consistently zero discrepancies in the dominant regime.
In a zero-sum combinatorial game setting, parity consideration can be central, determining whether you or the opponent gets the last move. Here, it remains a decisive factor in the following sense. The concept of a ‘last move’ is individualized; for all sufficiently large heaps (depending on $S$) both players have a last move. Sometimes you can sacrifice and secure some parity advantage, you win a last move that you otherwise would not have reached.
We are often interested in understanding play from the options of a heap of size $h$. This motivates us to denote the branches $H_{1}=h-s_{1}$ and $H_{2}=h-s_{2}$, and think of the $H_{i}$s as games with full game trees, rather than ‘heap sizes’.
@@
If a player sacrifices playing in PSPE from a heap of size $h$, independently of tie-breaking rules, they know that one of the two scenarios apply:
@@
More explicitly, consider (i). Here, their last move $s_{2}$ in $H_{1}$ is replcaed by $s_{1}$ in $H_{2}$, giving them a gain of $\Delta$ in $H_{2}$. Meanwhile, the first move in $H_{1}$ and $H_{2}$ is $s_{1}$ and $s_{2}$ respectively. Therefore, in the first move they have a loss of $\Delta$ in $H_{2}$. Thus, with the gain and loss balanced out in $H_{1}$ and $H_{2}$, both sacrifice and greedy play is in PSPE.
In case of (ii), the sacrificing player loses $\Delta=s_{1}-s_{2}$ in their first move. Since the slack, at the end of $H_{1}$, is strcitly less than $s_{2}$, the gain in the added last move in $H_{2}$ is strictly less than $s_{2}+\Delta=s_{1}$. Thus, the last move can only be $s_{2}$. Therefore, in the strictly dominant regime ($\Delta>s_{2}$), the $H_{2}$ branch gives a strcitly lesser utility and thus sacrifice cannot be in PSPE. In balanced regime ($\Delta=s_{2}$) a player performs equally in $H_{1}$ and $H_{2}$, and in non-dominant regime ($\Delta<s_{2}$), $H_{2}$ branch gives a strictly better utility. Thus, in both these cases sacrifice play is in PSPE.
Now, suppose a player sacrifices, but neither (i) nor (ii) applies. Then on the branch $H_{2}$ they cannot obtain an extra move and cannot upgrade a final $s_{2}$ to $s_{1}$. Thus, their subsequent utility from $H_{2}$ is the same as from $H_{1}$. But they are behind by $\Delta=s_{1}-s_{2}$ already after the first move. Hence the greedy branch $H_{1}$ gives strictly larger utility. This contradicts the assumption that the sacrifice is played in PSPE. Therefore in PSPE any sacrifice must be justified by either (i) or (ii). ∎
@@
Hence, concerning the two first items in the statement (so $X={\rm FvA}$), it suffices to prove that

### v2

# Tie-breaking in self interest cumulative subtraction games

@@ 1 Introduction
We study two-player alternating play perfect information games and we adapt the following definition of a self-interest symmetric cumulative subtraction game $S\subset\mathbb{N}=\{1,2,\ldots\}$. There is a single heap of pebbles, of size $h\in\mathbb{N}_{0}$, and the players alternately remove $s\in S$ pebbles from the heap, until the current player is not able to move from some position $h$ because, for all $s\in S,h-s<0$. The players are called Alice and Bob, and whatever they remove from the common heap, they put in their own individual pockets. Each player aims to maximize the size of their pocket (i.e., their cumulative utility) when the game ends.
@@
We adapt standard tools from game theory, particularly the classical backward induction procedure, to player profiles in Pure Subgame Perfect Equilibrium (PSPE). To avoid ambiguity in the recursive evaluation and to ensure the game tree is well-defined, we assume deterministic tie-breaking in cases of indifference. We refer to a pair of tie-breaking conventions, one for each player, as deterministic if they satisfy six properties. They cannot be random or hidden in any way, and they must be universal, meaning that they are applied to every position whenever there is indifference between options, with respect to PSPE-computation. Moreover, they must be uniform, meaning that the same conventions are used across all positions. They must be Markovian (or history-independent), that is, they cannot depend on previous actions in any way. Finally, they are unilateral, meaning that each player’s convention is independent of the other player’s choice. One can prove by induction that PSPE play is indeed well defined by using these tie-breaking axioms.
@@
Using a pair of deterministic tie-breaking conventions, players still maximize their own utility. However, when multiple options yield the same personal outcome, the choice between them may influence the opponent’s utility. We classify this influence as friendly when it benefits the opponent and antagonistic when it harms them. This effect is understood to be local, meaning that it depends only on outcome discrepancies within any given pair of deterministic tie-breaking conventions. As shown in the Appendix, page Appendix: Unilateral deviation for larger subtraction sets, such local choices can have globally counterintuitive consequences when viewed across ranges of possible conventions.
@@
Deterministic tie-breaking conventions are commitments that are applied consistently throughout the game tree, rather than as decisions made at individual nodes. For instance, given a pair of tie-breaking conventions, for every heap size, even one as large as $10^{10}$, a player is required to know whether they are indifferent. If so, for example, a player who adopts a friendly convention must then select a move that maximizes the opponent’s result, based on the already computed PSPE utilities of the resulting subpositions.
@@
Following standard assumptions in combinatorial game theory, we assume the existence of such perfectly rational and computationally capable players. When both players adopt the friendly convention, we refer to the resulting setting as FvF or the friendly convention. If both adopt antagonistic conventions, we denote it AvA or the antagonistic convention. These will serve as our primary self-interest frameworks throughout the paper. [^1] In our model, the tie-breaking rule is fixed in advance and is not selected as part of the PSPE; analyzing strategic choice over tie-breaking behavior would require a different model, such as a simultaneous-play normal form with mixed strategies. Here, we focus on the recursive structure and properties of combinatorial games instead.
@@
1.1 The symmetric outcome function
Consider an antagonistic self-interest cumulative subtraction game. Following [15], let us define the self-interest outcome function, under symmetric tie-breaking, of a game $S\subset\mathbb{N}$. The outcome for the two players is the ordered pair $o_{\scriptscriptstyle{\rm AvA}}(h)=(o^{1}_{\scriptscriptstyle{\rm AvA}}(h),o^{2}_{\scriptscriptstyle{\rm AvA}}(h))$, where $h\in\mathbb{N}_{0}=\mathbb{N}\setminus\{0\}$ represents the current heap size. The cumulative utilities (a.k.a individual outcomes) can be recursively calculated using the outcome of heaps of smaller size. Let the first player choose $s\in S$ as their first move in PSPE play. Then $o^{1}(h)=s+o^{2}(h-s)$ and $o^{2}(h)=o^{1}(h-s)$. Thus, after the first move, the heap size is reduced to hsh-s, and the order of the players is switched (since the next move is made by the originally second player). Here, whenever there is indifference (i.e., more than one action leading to the same PSPE cumulation for the current player), the action chosen minimizes the PSPE cumulation for the other player, representing antagonistic behavior. In the case of friendly cumulative play, the individual outcomes $o^{1}_{\scriptscriptstyle{\rm FvF}}$ and $o^{2}_{\scriptscriptstyle{\rm FvF}}$ are analogous, but the action chosen maximizes the other player’s cumulation in case of indifference, representing friendly behavior.
Note that these two settings remain symmetric, in the sense that the outcome pairs do not depend on “who starts”, Alice or Bob, and therefore, unless otherwise stated, we use the convention that Alice starts. Thus, the setting represents a distant relative of the classical ‘impartial’ combinatorial games. The outcome definition gives us an algorithm to compute perfect player results/utilities.
Let the discrepancy of player $i$ be $\delta^{i}(h)=o^{i}_{\scriptscriptstyle{\rm AvA}}(h)-o^{i}_{\scriptscriptstyle{\rm FvF}}(h)$, for $i\in\{1,2\}$. A valid question is whether this discrepancy will be non-zero. In Table 1 we list some initial outcomes and discrepancies for the two-action ruleset $S=\{3,5\}$.
@@ 1.1 The symmetric outcome function
$\delta^{2}(x)$ 0 0 0 0 0 0 0 0 0 0 0 0 0 0 $\boldsymbol{-1}$ 0
At heap size $14$, Bob loses one unit in the antagonistic setting. The PSPE move sequence in the friendly setting is 3-3-5-3, where Alice ‘sacrifices’ in her first removal. In the antagonistic setting, she instead plays ‘greedily’, and the PSPE sequence becomes 5-5-3, where Alice gets both the first and the last move. We experimentally discovered a number of such examples of discrepancies in the two and three-action settings, summarized in Section 4, with related conjectures.
The main result of this paper, Theorem 3, is that, if the subtraction set has size 2, then, for any heap size, each player’s utility is no worse in the FvF setting than in the AvA setting. We conjecture that this result continues to hold if $|S|>2$; see Section 2.
Remark 1.
Readers familiar with zero-sum games may note that the move sequence discussed in the context of Table 1 cannot appear in an optimal-play zero-sum setting, where player removals add and subtract to a common score, respectively. Alice would instead start by removing 5, and hence Bob must take 5, and Alice gets the last move by taking 3. This, of course, leads to a loss in “social welfare” as the whole heap would not be consumed. A related topic concerns similarities between the $\rm AvA$} antagonistic and the zero-sum settings. We explore discrepancies between these games in a follow-up paper [12].
If the players play several rounds of the same game, and both have so far both been committed to the friendly setting, could either player benefit by instead deviating to deterministic antagonistic?
The Penguin narrative will be revisited in the Appendix.
@@
Could any personalized player benefit from a unilateral deviation from a deterministic friendly tie-breaking?
This question requires a generalized outcome function, and we devote Section 2 to this.
1.2 Outline
Let us outline the content of this paper. In Section 2 we generalize the outcome function to four deterministic tie-breaking pairs and explain the road map to the main theorem for two-action rulesets. In Section 3, we prove the main theorem, Theorem 3, via a few lemmas, highlighting first player advantage, greedy play vs. sacrifice, tie-breaking monotonicity and more. In Section 4, we outline a follow-up project with several structurally detailed conjectures concerning FvF vs AvA (those are the conjectures that motivated this study), while in Section 5, we outline some conjectures by instead comparing AvA with zero-sum play. In Section 6, we discuss further open problems, and in Section 7, we mention some related work. Finally, in the Appendix (page Appendix: Unilateral deviation for larger subtraction sets), we illustrate how cases where $|S|>2$ can differ from two action rulesets, in globally counter intuitive ways.
2 A generalized outcome and the main result
In combinatorial games, we define perfect play outcome functions as backward induction algorithms to compute PSPE; in asymmetric cases, such as “partizan play”, perfect play result can be sensitive to who starts. Observe that in our setting, Question 1 induces two more settings, and a more general outcome function. Let us list all settings.
@@ 2 A generalized outcome and the main result
    Alice, as a starting player, acts friendly, and Bob is antagonistic, in case of indifference;
@@
    Bob, as the second player, acts friendly, whereas Alice is antagonistic, in case of indifference;
@@
This includes assymetric settings, by their tie-breaking commitments, although the ruleset itself is symmetric. However, by defining the notion of a dual setting, we can still use a relatively compact definition of “outcome”.
Let $\tau=\{{\rm FvF},{\rm AvF},{\rm FvA},{\rm AvA}\}$, and let $X\in\tau$. The dual of $X$, denoted $\mathsf{d}(X)$, is defined by:
@@
Given a tiebreaking convention, we usually refer to a pair of (EGT) utilities in PSPE simply as a (CGT) outcome. As mentioned, even if the ruleset is symmetric, since the players may have different tie-breaking rules, a game played is not necessarily symmetric in the strict sense. If one player is antagonistic at some stage of play, then they are committed to being antagonistic at all stages of play, and similarly for “friendly”. This is a situation that cannot appear in standard CGT impartial games. In this sense, the move options depend on their committed tie-breaking rule. To the best of our knowledge, such situations have not yet been studied in the literature.
Thus, there are four deviation cases to consider, listed as FvF $\rightarrow$ AvF, FvF $\rightarrow$ FvA, AvA $\rightarrow$ FvA and AvA $\rightarrow$ AvF. For example, in FvF $\rightarrow$ AvF, in the FvF setting, both players resolve ties in a friendly manner, while in AvF, Alice (playing first) deviates with an antagonistic tie-breaking rule.
In the case $|S|=2$, we here prove that a unilateral deviation from friendly tie-breaking to antagonistic ditto never benefits either player; as we will see, whenever starting from FvF, the deviating player will be unaffected, while the other player may suffer in some cases (Lemma 8). As a consequence, we will get:
@@
Consider any subtraction set $S$, with $|S|=2$, and suppose that both players act friendly in case of indifference. Then each player’s PSPE utility is never worse than if both players have antagonistic tie-breaking rules.
Our methods to prove this result do not apply in general if $|S|>2$, and we provide a counter example, when unilateral deviation is non-monotone, namely if $S=\{4,5,9\}$, in Table 3 in the Appendix. We will study the two unilateral deviations from FvF, to FvA and AvF, while the converse/symmetric deviations from AvA, to AvF and FvA are analogous. A unilateral deviation from friendly to antagonistic may, in rare cases, be beneficial to the deviating player; still, we have numerical evidence that Theorem 3 continues to hold for any subtraction set of size greater than 2 (see Section 6). However, we currently have no formal method to jump directly from both friendly to both antagonistic.
@@
Consider ordered pairs of integers $(a,b)$ and $(c,d)$. Then $(a,b)\leq(c,d)$ if $a\leq c$ and $b\leq d$, and similarly with $\geq$”.
3 Unilateral tie-breaking deviation and a proof of the main result
Suppose that both players follow deterministic tie-breaking conventions. We are interested in comparing the outcome under one such setting, where both players are universally friendly, with the setting in which exactly one of the players deviates, by committing to an antagonistic tie-breaking rule.
Whenever the subtraction set, $S=\{s_{2},s_{1}\}$, has exactly two actions, then we claim that the deviating player’s PSPE-utilities will be exactly the same as if they had remained friendly, and the other player’s PSPE-utilities cannot improve. This claim holds independently of who starts, and we prove it in Lemma 8 in this section.
In the following sections, we will fix the convention $s_{2}<s_{1}$.
Through Lemma 8, we see how the claims in the first paragraph combine to form an induction proof. For any two-action subtraction set $S$, let $\Delta=s_{1}-s_{2}>0$. A greedy move from $h\geq s_{1}$ is to play to $h-s_{1}$, and otherwise a player sacrifices. The endgame refers to the play to a heap size $h<s_{1}$. The slack is a leftover that cannot be attained by either player, and therefore refers to a heap of size $h<s_{2}$.
There is a general proof of the next lemma in the zero-sum setting [5]. Here, it only holds in the two-action setting. See Table 3 in the Appendix on page Appendix: Unilateral deviation for larger subtraction sets, where $S=\{4,5,9\}$ and where heaps of sizes $61,75,\ldots$ have strictly better PSPE utility for the second player.
@@ 3 Unilateral tie-breaking deviation and a proof of the main result
Recall $s_{1}>s_{2}$. Suppose that Alice plays greedily until she cannot play $s_{1}$. Then she gets utility $ks_{1}+\ell s_{2}$, some $k>0$, some $\ell\geq 0$. If Bob gets to play $k$ copies of $s_{1}$, then Alice has the first player advantage at the endgame too, and hence Bob cannot get more than $ks_{1}+\ell s_{2}$. If Bob can only collect k1k-1 copies of $s_{1}$, then he can at most get one more copy of $s_{2}$ than Alice, but their utility difference is at least $\Delta>0$ in Alice’s favor. ∎
@@
The usefulness of a sacrifice may appear limited while playing on a single heap of tokens. However, simple arithmetic shows that if $S=\{2,3\}$ then Alice benefits strictly when playing from a heap of size $7$, and if $S=\{3,4\}$, then she can benefit strictly by sacrificing twice, playing from $h=17$. There are many more such examples, which are analogous to the zero-sum setting studied in [5]. In two-action zero-sum games, the usefulness of a sacrifice is limited to basic arithmetic considerations for small heap sizes. However, self-interest play appears to be more multifaceted. As shown in Table 1, there are situations where, under a deterministic tie-breaking rule, Alice sacrifices solely to benefit Bob, and Bob’s immediate response is also a sacrifice. Note that the example occurs in the non-dominant regime. By contrast, the dominant regime generally exhibits less complex behaviour.
@@
The proof will use a strategy-stealing argument to establish a contradiction. Suppose that Alice sacrifices in her first move and profits strictly in PSPE play. This means that she must get at least $k$ more $s_{2}$ actions to compensate for her initial sacrifice, if $s_{1}=ks_{2}+r$, $0\leq r<k$. However, since $k\geq 2$, this requires that Bob also plays a sacrifice at some point. Otherwise, Bob gets at least $3\Delta-s_{1}\geq\Delta>0$ larger utility than Alice, while Lemma 5 assures that Alice, as a starting player, never gets worse than Bob in PSPE play. But Bob can replace this sacrifice with $s_{1}$, by stealing $ks_{2}+r\leq(k+1)s_{2}$ of Alice’s required removal segment. Hence, Alice’s first sacrifice cannot lead to a strict gain. ∎
A consequence of this result is that, in the case of dominance, $s_{1}$ is never a bad choice for the current player (if feasible); see Proposition 10.
In a zero-sum combinatorial game setting, parity consideration can be central, determining whether you or the opponent gets the last move. Here, it remains a decisive factor in the following sense. The concept of a ‘last move’ is individualized; for all sufficiently large heaps (depending on $S$) both players have a last move. Sometimes you can sacrifice and secure some parity advantage; you win a last move that you otherwise would not have reached.
We are often interested in understanding the play from the options of a heap of size $h$. This motivates us to denote the branches $H_{1}=h-s_{1}$ and $H_{2}=h-s_{2}$, and think of the $H_{i}$s as games with full game trees, rather than ‘heap sizes’.
@@
If a player sacrifices playing in PSPE from a heap of size $h$, independently of tie-breaking rules, they know that one of the two scenarios applies:
@@
More explicitly, consider (i). Here, their last move $s_{2}$ in $H_{1}$ is replaced by $s_{1}$ in $H_{2}$, giving them a gain of $\Delta$ in $H_{2}$. Meanwhile, the first move in $H_{1}$ and $H_{2}$ is $s_{1}$ and $s_{2}$ respectively. Therefore, in the first move, they have a loss of $\Delta$ in $H_{2}$. Thus, with the gain and loss balanced out in $H_{1}$ and $H_{2}$, both sacrifice and greedy play are in PSPE.
In case of (ii), the sacrificing player loses $\Delta=s_{1}-s_{2}$ in their first move. Since the slack, at the end of $H_{1}$, is strictly less than $s_{2}$, the gain in the added last move in $H_{2}$ is strictly less than $s_{2}+\Delta=s_{1}$. Thus, the last move can only be $s_{2}$. Therefore, in the strictly dominant regime ($\Delta>s_{2}$), the $H_{2}$ branch gives a strictly lesser utility, and thus, sacrifice cannot be in PSPE. In a balanced regime ($\Delta=s_{2}$), a player performs equally in $H_{1}$ and $H_{2}$, and in a non-dominant regime ($\Delta<s_{2}$), the $H_{2}$ branch gives a strictly better utility. Thus, in both these cases, sacrifice play is in PSPE.
Now, suppose a player sacrifices, but neither (i) nor (ii) applies. Then on the branch $H_{2}$ they cannot obtain an extra move and cannot upgrade a final $s_{2}$ to $s_{1}$. Thus, their subsequent utility from $H_{2}$ is the same as from $H_{1}$. But they are behind by $\Delta=s_{1}-s_{2}$ already after the first move. Hence, the greedy branch $H_{1}$ gives a strictly larger utility. This contradicts the assumption that the sacrifice is played in PSPE. Therefore, in PSPE any sacrifice must be justified by either (i) or (ii). ∎
@@
Hence, with regard to the first two items in the statement (so $X={\rm FvA}$), it suffices to prove that